The following is a set of data from a sample of n=11 items. Complete parts (a) through.(c). X   4   19    17   15   12   7   11   6   10   13   18 Y   8   38   34   30   24   14   22   12   20   26   36 Compute the covariance. Compute the coefficient of correlation.     How strong is the relationship between X and​ Y? Explain.                                                                                  Compute the sample mean for X, rounding to six decimal places. Upper X overbarXequals=37.63636437.636364 Compute the sample mean for Y, rounding to six decimal places. Upper Y overbarYequals=12.54545512.545455 Compute the covariance for X and Y, rounding to three decimal places. i need to know each step in finding the covariance.
Answer #1

0 121 ) - 21 - 3 - [12) mean ý = (39 Covariance ( (11) (1-) (V3) 8 n - 1 7 (y - 1234 -+ (-12)(2-29) -*(4-21) 11 - 1 - 50-0000(b) X values > x= 132 mean =12 [(x-Mx)}= SSx = 250 y values s = 964 mean = 94 2 Cy-My)2= ssy = 1000 x X and Y Combined hall E(C) X and y have because all boint a positive slope. a perfect positive correlation fall on stright line with